Find the slope of the line 3x-2y=12
Express your answer as a numeric value, using decimal notation.

Answer by stanbon(75887)

Find the slope of the line 3x-2y=12
-----
Solve for "y"::
2y = 3x-12
-----
y = (3/2)x -6
-----
Slope = (3/2)
====================
Cheers,
Stan H.
==================